Given: 9.89 glucose
Find: Li (g)
1
Expert's answer
2021-09-28T07:38:32-0400

Molar mass of Li = 6.941

Molar mass of glucose = 180.156

9.89/180.156 = 0.055mol

= 6.941×0.055 = 0.38g